OpenStudy (rayford):

What is the 22nd term of the arithmetic sequence where a1 = 8 and a9 = 56 ?

OpenStudy (raden):

a9 = a+8b, right ? with b is difference 1st term and the next term solve for b

OpenStudy (rayford):

ok i'm with you

OpenStudy (raden):

what's do u get for b ?

OpenStudy (rayford):

142?

OpenStudy (raden):

nopes we know that the generally form to find term of the arithmetic sequence is an = a+(n-1)b so if given a9=56 then a1+(9-1)b = 56 8+8b=56, what's value of b here ?

OpenStudy (raden):

an = a1+(n-1)b

OpenStudy (rayford):

150!

OpenStudy (raden):

8+8b=56 8b=56-8 8b=48 b=48/8=6

OpenStudy (rayford):

158

OpenStudy (raden):

next to find the 22nd term, subtitite value a1=8, b=6, n=22 to formula above an = a1+(n-1)b a22 = 8+(22-1)6 a22 = 8+21(6) a22 = 8+126 a22 = ?

OpenStudy (rayford):

134

OpenStudy (raden):

that's right

OpenStudy (raden):

for exercise for u, find the value of a50nd ?